OpenStudy (studygurl14):

Hello! Welcome to StudyGurl14's Study Session! Feel free to ask any questions and I'll try my best to walk you through it and give you an answer. :) Satisfaction guaranteed! (Maybe...lol) Medals, fanning, and/or testimonies are appreciated if you receive the answer you are looking for, but not in any way required. Thanks and let the questions begin! ~StudyGurl14

OpenStudy (rianiscool):

What is the coordinate of point T? A. 12.4 B. 12.5 C. 12.6 D. 12.7

OpenStudy (rianiscool):

c

OpenStudy (studygurl14):

Yep, C is correct. Good job

OpenStudy (rianiscool):

i am now a FAN OF YOU

OpenStudy (rianiscool):

Simplify the expression. 68 – 18 ÷ 3 • 6

OpenStudy (rianiscool):

68 – 18 ÷ 3 • 6

OpenStudy (rianiscool):

what isa the ?? for

OpenStudy (studygurl14):

Just use refresh button it will get rid of those

OpenStudy (rianiscool):

ok thank you

OpenStudy (studygurl14):

Using order of operations, you do the divide and multiplication first. And go from left to right. So... 18 ÷ 3 = 6 6 • 6 = 36 68 - 36 = 32 So 32

OpenStudy (anonymous):

lol

OpenStudy (rianiscool):

thank you but @kitkatNat1316

OpenStudy (rianiscool):

helped me

OpenStudy (rianiscool):

thank you

OpenStudy (rianiscool):

need help on another one

OpenStudy (anonymous):

@rianiscool glad to help! (:

OpenStudy (rianiscool):

i will just ask

OpenStudy (anonymous):

Alright (:

OpenStudy (anonymous):

lol

OpenStudy (rianiscool):

llllllloooooolllll

OpenStudy (anonymous):

What is your question @rianiscool (:

OpenStudy (rianiscool):

Evaluate. 7 × [8 + (16 – 9)] ÷ 5

OpenStudy (rianiscool):

i am try ing to get the hang of this so i am going to get some extra help

OpenStudy (anonymous):

Ok, remember PEDMAS? We are going to need this again!

OpenStudy (rianiscool):

ok

OpenStudy (anonymous):

@StudyGurl14 do you want to do this or do you want me too? :)

OpenStudy (studygurl14):

order of operations. do 16-9 first 16 - 9 = 7 Then do 8 + 7, which = 15 Now you have 7 × (15) ÷ 5 Now, do left to right. 7 × 15 = 105 105 ÷ 5 = 21

OpenStudy (rianiscool):

k

OpenStudy (studygurl14):

PEMDAS: Parenthesis Exponents multiplication and division addition and subtraction

OpenStudy (rianiscool):

please excuse my dear ant sally

OpenStudy (anonymous):

@rianiscool exactly!

OpenStudy (rianiscool):

;;;;;;;

OpenStudy (anonymous):

Anymore questions, @rianiscool ? (:

OpenStudy (rianiscool):

21 is the answer

OpenStudy (anonymous):

@rianiscool yep, good job!

OpenStudy (rianiscool):

ok

OpenStudy (anonymous):

If you need help with anything else, feel free to tag me!

OpenStudy (studygurl14):

same here

OpenStudy (rianiscool):

ok i will

OpenStudy (anonymous):

:) We are always glad to help!

OpenStudy (rianiscool):

when is the best time to tag you guys

OpenStudy (studygurl14):

I'm on and off all day long

OpenStudy (anonymous):

Anytime, but right now works for me!

OpenStudy (rianiscool):

ok